Answer 1 of 1 a. For Business segments Business Segment Unlevered Beta adjusted for cash Media Networks Parks & Resorts Studio Entertainment 1.0993 Consumer products 0.6752 Interactive Unlevered beta 1.002425 1.0313 6677 .0668 6034 1.0085 0.7024 1.2187 BUL = ßL/ (1 + (1-tarrate)( Debt/ Equity)) Beta unlevered adjusted for cash Ladjustedforcash BUL/(1- (Cash/Firmvalue) b. Unlevered betas for a firm Median regression beta for interactive business 1.03 (beta levered) DE for Disney 13% 0, 13 Marginal tax rate-36% Beta unlevered 0.95088 using BuL = ßL/ (1 + (1-tan-rate)(Debt/ Equity)) Adjusting for cash, Beta unlevered - 0.95088/(1- 0.1725) 1.1490997 C. Levered beta for the firm B1-BUL 1 (1- tarrate) (Debt/Equity)) Beta levered for Disney = 1.2445968Done_ hw3 problem2_data (2 of 6) Estimate Cost of Equity and Cost of Capital - Disney Step 1: Estimate Cost of Equity for Each Business Segment Comparable Firms Disney Unlevered BetaDisney Corrected for Cash Ratio (%) DE Corporate 7 Equity RiskCost of PremiumEquit Business Segments Comparable HS Rate Levered Beta Risk Free Rate Radio and TV Broadcasting Companies Theme Park & Entertainment Firms Movie Companies Media Networks 10.03% 36.10% 2.75% 5.76% Parks & Resorts Studio Entertainment 11.41% 20.71 % 36.10% 36.10% 2.75% 2.75% 5.76% 5.76% Consumer Products Interactive Dis Toy and Apparel Retailers; Entertainment Software Gaming Firms 117.11% 4 1.07% 13.10% 36.10% 36.10% 36, 1090 2.75% 2.75% 2.75% 5.76% 5.76% 5.76% Step 2: Estimate Cost of Capital for Each Business Segment Disney Disney Pre- Disney Afier- Corporate Ta Tax Cost of Tax Cost ofDebt CapitaCost of Capital Cost of Equity Rate Debt Debt Ratio 9.12% 10.24% Busimess segmentS Comparable HS Radio and TV Broadcasting Companies Theme Park & Entertainment Firms Movie Companies Media Networks 36, 10% Parks & Resorts 36. I 0% 36, 10% Studio Entertainment 17.16% Consumer Products Interactive Disney Toy and Apparel Retailers; Entertainment Software Gaming Firms 36. I 0% 36, 10% 36. 10% 53.94% 29. I 1% 1 1.58% ations Source: Applied Corporate Finance, Damodaran, 4th Edition.Problem 2: Estimate Cost of Equity_and Cost of Capital for Disney. Based on the data provided in Homework3_Problem2_Disney Cost of Capital.xls for Disney, calculate cost of equity and cost of capital for each business segment of Disney: a) Based on unlevered beta corrected for cash for each business segment from Problem 1, estimate levered beta for each business segment; free rate of 2.75%, and equity risk-premium of 5.76%); pre-tax cost of debt assuming a risk-free rate of 2.75%? What is its b) Estimated cost of equity for each business segment (assuming a risk- c) Estimate after-tax cost of debt assuming a cooperate tax rate of 36.1%? d) Estimate cost of capital for each business segment of Disney.
Answer #1

Please note:

  1. A screenshot of my excel file in show formula mode - so that you can see the formula in each cell and understand the mathematics.
  2. The row below the heading also gives you an idea as how each column has been calculated
  3. I have then attached my final table so that you can see the final answer after typing in the excel formula
  4. Part (a) and (b) have been calculated in first part of the table
  5. Part (c): is shown below
    1. Rating: A; Risk spread corresponding to rating A = 1%
    2. Risk free rate = 2.75%
    3. Pre tax cost of debt, Kd = Risk free rate + Risk spread = 2.75% + 1.00% = 3.75%
    4. After tax cost of debt = Kd x (1 - Tax rate) = 3.7%% x (1 - 36.1%) = 2.3963%
    5. I have used 3.7% and 2.3963% in second table. You now understand how they had been calculated.
  6. Part (d) is calculated in second table
  7. I have expressed all the results upto four places of decimal. Please round off the output as per your need.

Disney Part (b) Cost of equity Comparable Disney D/E Corporate Comparable firms Firms Risk Free rate Equity risk premium ltax rate tparta) 1ewsed beta unlevered beta |Ratio(%) 2 Business Segment Rf Bu D984) 0.0275 10.361-985.(1+(1F985)Dasǐ 0.0275 D87)0.0275 Cot otd 98611 E986)*D986) 0.0275 -C988 (1+1-E988) 0.0576 0.0576 0.0576 0.0576 0.0576 0.10030.361 984 Media Networks 985 Parks & Resorts 86 Studio Entertainment 987 Consumer Products 988 Interactive 89 Disney Operations 1.0313 024 1.0993 6752 1.2187 0.1141 G986+F 986, H986 0.361 0.361 0.361 2071 6987+5987 H987 G9aa+F988*H988 LH1-E98 988*(111-E988) 0.0275 0.4107 0.131 361 990 991 Step 2 Disney Pre Comparable Corporate tax tax Cost of Disney After tax Cost of 992 Business Segment Firms C Cost of equity rate debt debt Debt/ Capital Ratio Part (d) Cost of Capital Kd x(1-T) / Cx Kdx (1-T+(1-D/C)x Ke 993 994 Media Networks 995 Parks & Resorts 996 Studio Entertainment 997 Consumer Products 998 Interactive 99 Disney Operations 0.361 0.361 0.361 361 0.361 0.361 0375 -994 (1-D994) 0375 5) 0.0375-E996*11-D996) 0.0375 997(1-D997) 0375-E998*11-D998) 0.0912 0.1024 0.1716 0.5394 0.2911 0.1158 G994 F994+(1-G994)*C994 -G995*F995+(1-6995) C995 996+1 G996) C996 97 F997+( 1-G997). С997 998 F998+(1-6998) *C998 1984 985 G996*F -1987 =1988

Step 1
Business Segment Comparable Firms Comparable firm's unlevered beta corrected for cash Disney D/E Ratio (%) Disney Corporate tax rate (Part a) Levered beta Risk Free rate Equity risk premium Part (b) Cost of equity
Bu D/E T BL = Bu x (1 + (1-T) x D/E) Rf Rp Ke = Rf + BL x Rp
Media Networks 1.0313 10% 36%              1.0974 2.75% 5.76% 9.0710%
Parks & Resorts 0.7024 11% 36%              0.7536 2.75% 5.76% 7.0908%
Studio Entertainment 1.0993 21% 36%              1.2448 2.75% 5.76% 9.9199%
Consumer Products 0.6752 117% 36%              1.1805 2.75% 5.76% 9.5495%
Interactive 1.2187 41% 36%              1.5385 2.75% 5.76% 11.6119%
Disney Operations 13% 36%
Step 2
Business Segment Comparable Firms Cost of equity Disney Corporate tax rate Disney Pre tax Cost of debt Disney After tax Cost of debt Debt / Capital Ratio Part (d) Cost of Capital
Ke T Kd Kd x (1 - T) D / C D / C x Kd x (1-T) + (1-D/C) x Ke
Media Networks 9.07% 36.10% 3.75% 2.3963% 9.1200% 8.4623%
Parks & Resorts 7.09% 36.10% 3.75% 2.3963% 10.2400% 6.6101%
Studio Entertainment 9.92% 36.10% 3.75% 2.3963% 17.1600% 8.6289%
Consumer Products 9.55% 36.10% 3.75% 2.3963% 53.9400% 5.6911%
Interactive 11.61% 36.10% 3.75% 2.3963% 29.1100% 8.9293%
Disney Operations 36.10% 11.5800%